A Complete Bibliography of Publications in the Journal of Software Maintenance and Evolution: Research and

Practice

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

31 January 2018 Version 1.00

Title word cross-reference

mm [AHAD05].

12th [KTW09]. 15504 [PBPV10, SPPV10].

2009 [Rou10a].

978-0470-14707-8 [Cha09]. 981-238-426-X [Sea08].

ability [OEMD08]. Abran [Cha09]. abstracting [JHHF08]. abstraction [ZR07]. abstractions [BLE^{+}08]. access [Kos02]. accurate [ZR07]. activities [ABRF11, JK09]. activity [WWS07]. adaptive [BK03]. adopt [HMG03]. adopting [TGH^{+}04]. adoption [MM11b]. Advances [Sea03]. affect [RDT^{+}07]. affecting [Hea02]. affects [BKS07]. Ageing [Vis01]. Agile [TWFF11, BDJ^{+}11, FGSO10, HMG03, MM11b, OELC06, OEMD08]. Alain [Cha09]. Alamitos [Cha09]. algorithm [DST07]. alignment [KG11, TRPG06]. among [KG11]. analysing [TV01]. Analysis [ERS04, ZHLG08, AP05, BKS07, DM07, KG06, MMD03, MdO11, NMS11, NGA10, RABT08, SZ01, SCRS08, UKI01, WBHS08, ZLL^{+}10, ZYXX02]. analyze [Del09, GD06]. Analyzing [GDG06, MRG04, PVvV05, BAA09]. annotated [DGR^{+}11]. Announcement [CC01]. APIs [DJ06]. application [BK03, GCT01, KNTRS08, WWS07, ZR07]. Applications [Ger09, DST07, DFT04, GJ09, RDT^{+}07, TR04, ZGFH09]. applying [WBHS08]. approach [Bal04, BSGK10, BG02, Dan10, DFT04, HBR10, HKI08, JHHF08, LYC01, Lit04, NM05, PGRP07, RHH09, RT11, RSSX11, SK03, SCRS08, TK04, TK01, WSWM06, ZLL^{+}10]. approaches [KCM07]. April [Cha09]. Architectural [BKP05, CDN10, GZ02, ZYXX02]. Architecture [CS03, ASR09, Del06, GCRFPL02, HE02, RSSX11]. Architecture-based [CS03]. architecture-centric [ASR09]. Armstrong [Sea08]. articles [Rou10a]. Assessing [BMP05, JK09, KKSM08, OELC06, Rou10b]. assessment [ASR09, AT04, HE10, Rou10b, SPPV10]. assets [Vee04]. assisted [SK03]. assurance [Li09]. atomicity [AP05]. attribute [NP07]. attribute-oriented [NP07]. Authored [Cha09, Sea08]. Automated [TZ03, JHHF08]. Automatic [TRPG06, ZD08, ZR07]. Automating [ACDS02, MNX05]. automation [BKS07]. automotive [MPÖ10]. AVal [NP07]. avoidance [GHD03]. aware [KR11].

Baan [vGBB05]. Bad [ZHB11]. Balancing [Dan10, GJ09]. Based [DAH08, ASKK03, AS05, BSGK10, CDTM10, CMJ04, CPU09, CS03, CMW02, DHT^{+}08, FG04, HKI08, LBLA04, LCC^{+}03, MTN11, NMS11, NT07, OBS04, OÓ08, PGRP07, PM04, RABT08, RPTG08, RSSX11, SCRS08, TK04, TV01, UKI01, VB02, WSWM06, ZLL^{+}10, ZDM10]. Bayes [Ste02]. BBN [BSGK10]. behavior [ERS04]. Behavioural [LKR02, NGA10]. belief [LYC01]. Best [MY10, HP11, Li09]. better [Don03]. biting [ZLL^{+}10]. biting-down [ZLL^{+}10]. blame [Qua10]. Book [Cha09, Sea01, Sea03, Sea06, Sea08]. bottom [OBS04]. bottom-up [OBS04]. BPM [BDJ^{+}11, EGH^{+}10, MY10, RMSL11]. building [ASP02]. Business [KKGL10, DGR^{+}11, HBR10, TGH^{+}04, WWEL10, ZGFH09].

C [BKP05, JMK01]. C-POSH [BKP05]. CA [Cha09]. call [ZR07]. Capabilities [RABT08]. Capabilities-based [RABT08]. capability [BMP05, McB10, Rou10a, Rou10b]. Capturing [HBR10]. Cartography [KELN10]. Case [PAK01, ACDS02, BK03, BSGK10, Dan10, Del06, ERS04, GZ02, JBBvN06, JL06, KMPY06, MPÖ10, MH05, MRE11, NFC01, OAW01, OEMD08, SMKN11, WSH01, vGBB05]. cases [AS05]. catalog [MdO11]. centric [ASR09]. certification [NS10]. challenges [BDJ^{+}11, TGH^{+}04]. chance [EGH^{+}10]. Change [ZYXX02, BMZ^{+}05, GJ09, GCS^{+}10, Hea02, KKSM08, Qua10, RABT08, RD10]. changes [BAA09]. Characteristics [NFC01]. Chartwell [KNTRS08]. checking [MM11a]. class [AS05, KKOL11, OEMD08, SJC02, SC06]. class-based [AS05]. classes [AS05, OEMD08, ZD08]. Classification [RT11, BG10, CHLW06]. classifier [Ste02]. client [Ng01]. clone [DNR06]. clones [HKI08, JL06, KG06, RC10]. clustering [DST07, RPTG08, SK03]. clusters [RD10]. CMMI [LWMW10, PBPV10]. CMMI-DEV [PBPV10]. cockpits [HM10]. Code [ZHB11, DM07, HKI08, LYC01, MKB09, MMD03, TK01, TZ03]. codifying [CDN10]. cohesion [MWM^{+}11]. collaborative [DGR^{+}11]. Combining [EGH^{+}10]. commerce [Tay03]. communication [Qua10]. company [LWMW10]. comparison [ED07]. competitive [DST07]. completeness [OYE07]. Complex [MM11b, Don03]. complexity [OEMD08, SC06]. Component [AP05, BLE^{+}08, DHT^{+}08, KMPY06, SK03, WSWM06]. component-based [WSWM06]. components [TV01]. composition [AP05]. Comprehend [Del09]. Comprehension [ZHLG08, OBS04, RDT^{+}07]. computation [AFGK02]. Computer [Cha09]. Computing [Bla01, EFP07]. Concept [Bol08, LYC01]. concept-oriented [LYC01]. conception [WWEL10]. Concepts [Sea01, Sea08]. concerns [HE02, MW02]. Conference [KTW09, RG06, vdBCG04]. configuration [AFGK02]. conformant [SPPV10]. consideration [LKP11]. considering [XXN11]. consistency [HE02]. Consistent [BG02, KELN10]. construction [WWEL10, ZR07]. context [AT04, Dan10, KCM07, MPÖ10]. Continuous [Cha09, Del06, KKGL10, MH05]. contradiction [EGH^{+}10]. contribution [Qua10]. controlled [CS03]. coordination [AFGK02]. Copyright [Cha09, Sea08]. Corporate [Sea06]. correction [NT07]. corrective [KM02]. cost [MPÖ10]. COTS [MRE11]. cover [Cha09]. coverage [Kan08, NMS11]. creational [NGA10]. CSMR [KTW09, vdBCG04]. cultural [TWFF11]. culture [Qua10, TWFF11]. current [ZHB11]. curves [SLSN10]. customization [HM10, JBBvN06, Lig01].

data [CDTM10, FG04, OYE07, Vis01]. data-intensive [Vis01]. database [BG02, PGRP07, RS05, TZ03]. databases [RS05]. dataflow [Li09]. de-localized [VB02]. debugging [CMJ04]. decision [ABRF11, Ng01]. decisions [CDN10]. decomposition [ZLL^{+}10]. deeper [Kan08]. defects [RHH09]. Defining [MdO11]. delivery [JBBvN06, Ste02]. departmentalization [YJ02]. dependence [Bal04]. dependencies [KM07, TZ03]. dependency [LKP11]. deployment [JBBvN06]. Design [vGBB05, BCG^{+}05, CDN10, KKOL11, NGA10, OELC06, RABT08, RDT^{+}07, Sch07, SH07, TK04, TA01, UKI01]. designing [BCG^{+}05]. designs [MAP09]. detection [DNR06]. Determinants [ZWP03]. determination [Rou10a]. DEV [PBPV10]. developer [JK09]. Developing [SMKN11]. Development [MH05, BSGK10, CD10, CJD11, ED07, FGSO10, JBBvN06, JK09, KG11, MM11b, PSTN06, TMSM02]. diagrams [BMS^{+}08, SC06]. diffusing [BCG^{+}05]. digital [PSTN06]. Dimensions [DT03]. directing [WWS07]. directives [Spi09]. diseconomies [CJD11]. Displaying [Bal04]. distributed [Ala11]. diverse [LCC^{+}03]. diversity [RSL03, SCSR03]. do [DJ06]. documenting [CDN10]. domain [LYC01]. down [ZLL^{+}10]. driven [FGSO10]. during [GHD03]. Dynamic [GJ09, ZHLG08, Li09, NGA10, WBHS08]. DynaQoS [Li09]. DynaQoS-RDF [Li09].

early [BMP05]. easing [GZvDH09]. eBusiness [AT04]. Economies [CJD11]. Editorials [BDDT07, CDR^{+}11b, CDR11a, DM07, DAH08, HZD10, JSB10, KTW09, MY10, RMSL11, Rou10a, ZHLG08]. effect [Bla01]. effective [RO03]. effectiveness [DNR06]. efficacy [JK09]. effort [ASKK03, BMP05, HE10, Li09, Yu06]. electronic [Tay03]. elicitation [TZ03]. embedded [CS03]. Emergent [HS01]. Empirical [NT07, BKS07, FSC^{+}11, LFB07, OÓ08, OEMD08, RABT08, RD10, RC10, YJ02]. Empirical-based [NT07]. enabling [BDJ^{+}11]. Encapsulating [BLE^{+}08]. end [KM04]. enforcing [BCG^{+}05]. engineered [BMS^{+}08]. engineering [CMW02, DFT04, HZD10, Kos03, Lit04, PGRP07, SC06]. enhance [HMG03]. enhancement [SPPV10]. enough [HP11]. enterprise [ACDS02, Ng01]. enterprises [LWMW10]. entropy [EGJ02]. environment [DT03, VB02]. era [CDR11a]. ERP [Dan10, HS01, JBBvN06, Lig01, NFC01, vGBB05]. error [NT07, OEMD08]. error-correction [NT07]. error-prone [OEMD08]. errors [AP05]. estimation [ASP02, ASKK03, Dan10, MPÖ10]. European [vdBCG04, NS10]. Evaluation [Cha09, LKP11, TV01, BMS^{+}08, Del06, KG11, Kos02, MKB09]. events [JHHF08]. evidence [Rou10b]. Evolution [BDDT07, BT04, CDR^{+}11b, CHLW06, KF05, MRG04, PVvV05, Sea06, AT04, BKS07, BAA09, CHK^{+}01, CS03, DL09, Del06, Del09, Don03, FG04, GHD03, Ger04, GD06, GZvDH09, GDG06, KCM07, Koc07, LBLA04, LKR02, LCC^{+}03, MH05, MW02, NM05, OYE07, PSTN06, RO03, RHH09, RS05, RAH03, WSWM06, ZYXX02]. evolutionary [KM07]. evolve [DJ06]. Evolving [GCRFPL02, MWM^{+}11, MNX05, Zen05]. examination [GCS^{+}10]. example [LKR02]. execution [JHHF08]. expanding [LWMW10]. Expectation [OBS04]. Expectation-based [OBS04]. Experience [Sch07, GZvDH09, JS02, RSSX11, WJRR08]. Experimental [Kos02]. experimentation [BMP05]. expertise [Ala11]. Exploratory [ZDM10, SEW^{+}06]. Extended [WJRR08]. extensible [NP07, Zen05]. extracting [TK01].

facilitating [JK09]. factors [BSAM06, Hea02]. Father [CDR^{+}11b]. fault [OAW01]. fault-proneness [OAW01]. feature [FG04, GDG06, SEW^{+}06]. features [BMS^{+}08, NT07]. Feed [MMD03]. Feed-forward [MMD03]. Field [ASP02, ZWP03]. file [Spi09]. Finding [SLSN10]. findings [FSC^{+}11]. firewall [WJRR08]. flexibility [GJ09]. focused [ERS04]. formal [Bol08]. Formalizing [MVDJ05]. formed [LFB07]. Fortran [EFP07]. forward [MMD03]. foundations [CHLW06]. fragments [TK01]. framework [AS05, DGR^{+}11, GCT01, KR11, Ng01]. front [KM04]. front-end [KM04]. Function [UKI01, ASKK03, RC10]. Function-point [UKI01]. functional [ASP02, HE02]. functionality [TK01]. future [GZvDH09].

generation [ED07, EFP07]. generative [JL06]. Girish [PC01]. Goal [HM10]. Goal-oriented [HM10]. grammar [PM04, WBHS08]. grammar-based [PM04]. Graph [CMW02, MVDJ05, ZR07]. Graph-based [CMW02]. graphs [Bal04]. grow [MKB09]. grow-and-prune [MKB09]. Grubb [Sea08]. Guaranteed [GHD03]. Guest [Rou10a]. GUI [MNX05].

handling [MRE11]. Harmonizing [PBPV10]. header [Spi09]. healthcare [MM11a, Qua10]. heterogeneous [MH05]. hierarchical [Bal04, ZLL^{+}10]. hierarchies [SJC02]. high [HW02]. highly [OEMD08]. history [GD06]. Hoboken [Cha09]. human [Ala11]. hypermedia [GCRFPL02]. Hypertext [KSP04, Kos02].

ICMM [Kol11]. ICSM2005 [RG06]. Identification [NGA10, MKB09, ZD08]. identifiers [LFB07]. Identifying [DST07, HW02, AT04, HKI08]. IEC [PBPV10, SPPV10]. IEEE [Cha09, RG06]. Impact [JS02, FGSO10, ZYXX02]. implement [TK01]. implications [Lig01]. improve [Del09]. improved [JMK01]. improvement [NS10, NT10, Rou10a, Cha09]. Improving [RABT08, RPTG08, TK04, WBHS08, BSGK10, ZDM10]. include [Spi09]. inclusion [TZ03]. inconsistency [GHD03]. Indicating [OYE07]. indicators [BMP05, MdO11]. Indirectly [Yu06]. individual [BAA09]. Industrial [SEW^{+}06, GZvDH09, RSSX11, SMKN11]. industry [BKP05]. Inference [TA01, OBS04, WBHS08]. inference-based [OBS04]. influence [BSAM06, YJ02]. informal [MMD03]. Information [Qua10, Sea06, ASR09, Hea02, KKOL11, KSP04, MMD03]. infrastructures [DT03]. inheritance [SJC02]. inhibitors [MPÖ10]. input [NT07]. inspection [ZDM10]. inspections [Kol11]. instability [OELC06]. Integrated [JBBvN06, KKOL11]. Integrating [LCC^{+}03, SPPV10, Ala11]. intensive [KG11, Vis01]. Interface [Ger09, AS05]. internal [BMP05]. International [RG06]. Internet [BK03]. Interscience [Cha09]. Interview [PC01]. Introduction [HZD10, MY10, ZHLG08]. Investigating [LWMW10]. investigation [Koc07, RD10]. ISBN [Cha09, Sea08]. ISO [PBPV10, SPPV10]. ISO/IEC [PBPV10, SPPV10]. Issue [BDDT07, DAH08, MY10, RMSL11, DM07, HZD10, KTW09, ZHLG08]. issues [LKR02, TMSM02, Tay03]. iterative [OELC06, OEMD08].

Java [CDTM10, GJ09, HKI08, NP07, ZR07]. Jersey [Sea08]. just [HP11].

KERIS [Zen05]. Key [BDJ^{+}11, ZD08]. keyword [RPTG08]. keyword-based [RPTG08]. knowledge [BCG^{+}05, Del09, LYC01, ZHB11].

label [BG10]. language [GZ02, UKI01]. languages [ED07]. Large [GCT01, MRG04, PVvV05, ACDS02, FG04, GZvDH09, Koc07]. Large-Scale [MRG04, PVvV05, FG04, Koc07]. Latitudinal [SCSR03]. layered [GCRFPL02]. layout [KELN10]. learning [ZDM10]. legacy [AT04, CDTM10, GZ02, HW02, KMPY06, LBLA04, RO03, Vee04, Vis01]. legal [Tay03]. Lehman [CDR^{+}11b]. levels [PBPV10, TWFF11]. library [ZR07]. line [KMPY06]. literature [BAA09]. lived [LKR02]. localization [KM07]. localized [VB02]. location [Bol08, SEW^{+}06]. logs [JHHF08]. long [LKR02]. long-lived [LKR02]. longitudinal [BKS07, SCSR03].

maintainability [ASR09, HMG03, KKSM08, SJC02]. maintainers [KSP04]. Maintaining [Tay03, RSSX11, VB02]. Maintenance [AHAD05, Cha09, DAH08, KTW09, RG06, Sea01, Sea03, Sea08, TMSM02, vdBCG04, ASP02, ASKK03, ACDS02, BAA09, BK03, BSAM06, Bla01, BMP05, CHK^{+}01, Don03, ED07, GCT01, HO05, HW02, HS01, JBBvN06, JS02, KM02, Kos03, Lig01, LKL08, LCC^{+}03, NFC01, Ng01, NT07, PAK01, SZ01, Sch07, UVG01, WSH01, WWS07, XXN11, YJ02, Yu06, Sea06]. Management [Cha09, JSB10, Sea03, ABRF11, ACDS02, KM02, KKOL11, KR11, Qua10]. manager [NS10]. Managing [Sea06]. manipulation [DM07]. Manny [CDR^{+}11b]. MANTOOL [UVG01]. many [DHT^{+}08]. many-to-one [DHT^{+}08]. market [FGSO10, JBBvN06]. market-driven [FGSO10]. MASS [KG11, JBBvN06]. mass-market [JBBvN06]. matching [DNR06]. Maturity [AHAD05, KM02, Kol11, PBPV10]. MDA [PGRP07]. MDA-based [PGRP07]. measurement [ASP02, McB10, NM05, SMKN11]. measures [WWS07]. Measuring [SC06]. Medi [CD10]. media [MH05]. medium [LWMW10]. medium-sized [LWMW10]. memory [CDR^{+}11b]. merging [HKI08]. meta [TK04]. meta-pattern [TK04]. method [JMK01, KG11, ZDM10]. methods [BG10, MM11b, TWFF11]. metric [EGJ02, HKI08, TK04]. metric-based [HKI08, TK04]. Metrics [SJC02, BMP05, MWM^{+}11, OEMD08, PM04, SLSN10]. Microsoft [Ger09]. Migrating [CDTM10, Lit04, GZ02]. Migration [Ger09]. mine [HMG03]. Mining [KM07, LKL08, BG10, KCM07, RT11]. Mismatch [MRE11]. miss [RC10]. mitigation [HO05]. Model [AHAD05, MM11a, ASKK03, Ala11, BSAM06, Hea02, Kol11, LCC^{+}03, MKB09, NMS11]. Modeling [ABRF11, GD06, RHH09, KKGL10]. Modelling [BLE^{+}08, LBLA04, OAW01, UKI01, LKR02]. models [ASP02, CPU09, HBR10, HP11]. modifiability [Vee04]. modification [FG04]. module [LKP11]. modules [Zen05]. motifs [NGA10]. much [HP11]. multi [LKP11, VB02]. multi-site [VB02]. multi-tasking [LKP11]. multilingual [TRPG06]. multiple [HP11, KKSM08, NFC01].

Naive [Ste02]. Naive-Bayes [Ste02]. Near [RC10]. Near-miss [RC10]. needs [KSP04]. network [Ala11]. networks [Del09, MMD03, NT10]. neural [MMD03]. NJ [Cha09]. non [GCS^{+}10]. non-reusable [GCS^{+}10]. notations [RDT^{+}07].

object [AS05, ED07, KR11, NMS11, OELC06, OEMD08, RS05, Sch07, TA01, ZLL^{+}10]. object-aware [KR11]. object-oriented [AS05, ED07, NMS11, OELC06, OEMD08, RS05, Sch07, TA01, ZLL^{+}10]. objective [Rou10b]. Observe [HMG03]. Observe-mine-adopt [HMG03]. OMA [HMG03]. one [DHT^{+}08]. online [WSWM06]. ontologies [MWM^{+}11]. ontology [OYE07]. OO [MTN11]. open [Koc07, RC10, Yu06, ZDM10]. open-source [Yu06]. opportunities [HKI08]. Optimal [XXN11]. Optimized [KNTRS08].

Optimizing [Spi09]. Organisational [McB10]. Organizational [PSTN06, SPPV10, TWFF11, YJ02]. oriented [AS05, ED07, HM10, LYC01, NMS11, NP07, OELC06, OEMD08, RS05, Sch07, TA01, ZLL^{+}10]. OSS [JK09]. outsourced [BSAM06]. overlay [CDTM10].

packaged [GCT01]. pages [Cha09, DST07, Sea08]. Pair [BCG^{+}05]. paperback [Sea08]. Papers [MY10]. paradigms [LCC^{+}03]. paradox [HP11]. Parametric [EFP07]. Parikh [PC01]. patches [RT11]. pattern [GZ02, TK04]. patterns [TA01]. PCODA [ZHLG08]. Peer [WWEL10]. Peer-to-peer [WWEL10]. Penny [Sea08]. performance [LKP11, Lit04, MdO11]. perspective [McB10, Ng01]. PHILharmonicFlows [KR11]. Piecemeal [GZ02]. planning [Ng01]. point [UKI01]. points [ASKK03]. Policy [SZ01, XXN11]. POSH [BKP05]. position [YJ02]. practice [BCG^{+}05, HP11, PAK01, RAH03, Sea08, Sea01]. predict [OEMD08]. Predicting [Ste02, Yu06]. Preparing [CDR11a]. preservation [vGBB05]. principles [FGSO10, WBHS08]. Problem [KM02, FG04, SEW^{+}06]. Problems [KM04]. Process [JSB10, RSL03, AHAD05, BKS07, BSGK10, BG10, Don03, ED07, HBR10, KKGL10, KR11, McB10, MdO11, NS10, NT10, Qua10, Rou10a, Rou10b, SCSR03, SPPV10, TMSM02, TV01, UVG01]. processes [Ala11, DGR^{+}11, LKR02, MM11a, RAH03, WWEL10, ZGFH09]. processing [PSTN06]. product [FGSO10, JBBvN06, KKOL11, KNTRS08, KG11, KMPY06, RAH03, WWS07, vGBB05]. profile [RSSX11]. profile-based [RSSX11]. profiles [GCS^{+}10]. program [EFP07, NMS11, ZHLG08]. programming [GJ09, JL06, NP07]. programs [JMK01, TK01]. project [ASKK03, Dan10, HE10, MM11b, Ste02]. projects [ABRF11, Don03, JK09, Koc07, OELC06]. prone [OEMD08]. proneness [OAW01]. proper [HO05]. properties [MM11a]. prototyping [BG02]. Provop [HBR10]. prune [MKB09].

QoS [Li09]. QoS-assurance [Li09]. qualitative [RAH03]. quality [HP11, OYE07, TK04].

rates [Ste02]. RDF [Li09]. re [CMW02, Kos03, PGRP07]. re-engineering [CMW02, Kos03, PGRP07]. reading [TV01]. real [PAK01]. real-time [PAK01]. Recommending [RD10]. reconfiguration [Li09]. reconstruct [Ger04]. reconstruction [DL09]. Recovering [ZGFH09]. recovery [LYC01, NT07]. recurrent [MMD03]. Reducing [FSC^{+}11]. reduction [CPU09]. Reengineering [KTW09, vdBCG04, DT03]. Refactoring [KMPY06, DJ06, HKI08, OÓ08]. refactorings [MVDJ05]. reflection [BKP05]. reflections [TWFF11]. Reflexion [BLE^{+}08]. Regression [CPU09, MAP09, JMK01, MTN11, MNX05, WJRR08]. Relating [HE02]. Relationship [Ala11]. relationships [HS01]. release [JBBvN06, KKOL11]. releases [KNTRS08, KKSM08, OAW01, vGBB05]. remedies [Vis01]. remotely [CS03]. repeated [WSH01]. report [FG04, GZvDH09, RSSX11, Sch07, WJRR08]. repositories [KM07, KCM07, SCRS08]. repository [GZvDH09]. requirements [ABRF11, AT04, CMJ04, CPU09, FSC^{+}11, HE02, KKOL11]. research [GCT01, Kos03]. resource [Ng01]. reusable [GCS^{+}10]. reuse [KMPY06, SPPV10]. Reusing [AS05]. Reverse [DFT04, BMS^{+}08, HZD10, Kos03, SC06]. reverse-engineered [BMS^{+}08]. Review [Cha09, Sea01, Sea03, Sea06, Sea08, BAA09, ZHB11]. revision [LYC01]. Revitalizing [Vee04]. ripple [Bla01]. risk [FSC^{+}11]. ROC [SLSN10]. role [LWMW10]. roles [HS01]. rules [LFB07, LKL08].

sampling [BG02]. Scale [MRG04, PVvV05, CJD11, FG04, Koc07]. scales [McB10]. SCAM [DM07]. scenario [ABRF11, TV01]. scenario-based [TV01]. scheme [CHLW06, NS10]. Scientific [Sea08, EFP07]. SDL [CPU09]. Search [DHT^{+}08, DAH08, OÓ08]. Search-Based [DAH08, DHT^{+}08, OÓ08]. segmentation [SCRS08]. segmentation-based [SCRS08]. selected [Rou10a]. selecting [JMK01]. selection [MRE11]. semantic [CS03, EGJ02]. semantically [DGR^{+}11]. Separating [AFGK02]. Separation [MW02, HE02]. sequence [BMS^{+}08, OAW01]. series [RHH09]. server [MH05]. services [Lit04, SH07, TGH^{+}04]. Seventh [vdBCG04]. signal [PSTN06]. similar [DST07]. similarity [MKB09]. simulation [ABRF11]. Site [BDDT07, BT04, RPTG08, TMSM02, VB02]. sites [ERS04, TRPG06]. size [ASP02, LWMW10]. sized [LWMW10]. skills [JS02]. SM [AHAD05]. small [LWMW10]. small- [LWMW10]. SME [BSGK10]. Smells [ZHB11]. Social [NT10, BDJ^{+}11, EGH^{+}10]. Society [Cha09]. soft [Cha09]. Software [CDR^{+}11b, DAH08, HE10, HO05, KF05, Koc07, KTW09, Kos03, KELN10, MPÖ10, MRG04, PVvV05, RG06, Rou10a, Sea03, vdBCG04, ASP02, ASKK03, ABRF11, ASR09, AHAD05, ACDS02, BKS07, BAA09, BSAM06, BSGK10, Bla01, BDJ^{+}11, BMZ^{+}05, BLE^{+}08, CHK^{+}01, CJD11, CHLW06, DL09, Del06, Del09, Don03, ED07, EGH^{+}10, FG04, FGSO10, GCT01, GHD03, GCRFPL02, Ger04, GD06, GZvDH09, GDG06, GCS^{+}10, HW02, HMG03, HE02, HM10, HKI08, JBBvN06, KCM07, KG06, KG11, KMPY06, Kol11, KSP04, KKSM08, LKP11, Lig01, LKL08, MAP09, MTN11, MNX05, MW02, MM11b, NFC01, NS10, NT10, NM05, OBS04, OEMD08, PSTN06, PAK01, PM04, RHH09, RT11, RSSX11, RD10, RC10, RAH03, SZ01, Sch07, SLSN10, SCRS08, UVG01, VB02, WBHS08, WSWM06, WWS07, XXN11, YJ02, Yu06, ZD08, Zen05]. software [ZWP03, ZDM10, vGBB05, AHAD05, Cha09, Sea01, Sea08]. Solutions [Sea03]. some [LKR02]. source [DM07, Koc07, LYC01, MMD03, RC10, TK01, TZ03, Yu06, ZDM10]. SPE [CHLW06]. Special [BDDT07, DM07, DAH08, KTW09, MY10, RMSL11, HZD10, ZHLG08]. specification [DGR^{+}11]. specifications [Bol08, UKI01]. SPICE [CD10, Rou10a]. Splitting [GZvDH09]. stability [OYE07]. Stable [MWM^{+}11]. staffing [KNTRS08]. stakeholders [KG11]. state [Hea02, RAH03]. static [AP05]. Statistical [TR04]. statistically [OAW01]. story [DJ06]. strategies [KG11]. strategy [AT04, CMJ04, RO03]. string [DNR06]. structure [ZLL^{+}10]. structures [CDTM10, Kos02]. Studies [Rou10b, ASP02]. study [ACDS02, BK03, BSGK10, Dan10, Del06, ERS04, GZ02, JBBvN06, JL06, KMPY06, LFB07, MPÖ10, MH05, MRE11, NFC01, OÓ08, OAW01, OEMD08, PAK01, RD10, RC10, SEW^{+}06, SMKN11, WSH01, YJ02, ZWP03, vGBB05]. subsequent [vGBB05]. subset [HO05]. substitution [DHT^{+}08]. suite [CPU09, PM04]. support [ABRF11, BKP05, CDN10, Don03, KM04, KSP04, MdO11, RD10, TRPG06, ZYXX02]. Supporting [KG06, UVG01]. survey [BMS^{+}08, FSC^{+}11, KCM07, Kos03, MAP09, RAH03]. symptoms [Vis01]. system [AT04, CPU09, Hea02, HKI08, LKP11, Vis01, ZD08]. systematic [Ala11]. Systems [Sea06, CS03, CHLW06, Don03, GCRFPL02, GCS^{+}10, KKOL11, KG06, LBLA04, Li09, RO03, SZ01, SMKN11, Tay03, VB02, ZLL^{+}10].

Tailoring [Don03]. Takang [Sea08]. targeted [BLE^{+}08]. tasking [LKP11]. tasks [BG10]. taxonomy [BMZ^{+}05, KCM07, RS05]. technical [TGH^{+}04]. technique [JL06, TV01]. techniques [ZD08]. Technologies [Sea03]. technology [Qua10, KNTRS08]. telecommunications [PAK01]. temporal [LKL08, SCRS08]. Test [NMS11, RAH03, AS05, CPU09, Kan08, SH07]. Testing [VB02, AS05, CMJ04, MAP09, MTN11, MNX05, SH07, TR04, WJRR08]. tests [JMK01]. text [RT11]. their [OEMD08]. thematic [KELN10]. theory [ASR09, KKOL11, McB10]. third [ED07]. third-generation [ED07]. threshold [SLSN10]. throughout [OYE07]. time [PAK01, RHH09, XXN11]. timed [MM11a]. tolerance [RABT08]. tool [BMS^{+}08, CDN10, SH07, UVG01]. tooling [WWEL10]. tools [CMW02, SEW^{+}06, VB02]. trails [Ger04]. training [NS10]. transactions [TZ03]. transformations [MVDJ05, TK04]. transparency [GJ09]. Types [CHK^{+}01].

U.S.A [Cha09]. UML [MAP09, MTN11]. UML-based [MTN11]. Unanticipated [KF05]. unavailable [XXN11]. uncertainty [Dan10]. Understanding [BAA09, BMS^{+}08, Kan08, RPTG08]. unified [LCC^{+}03, UKI01]. Unifying [JL06]. update [AP05, GJ09]. upgrade [Ng01, SZ01]. US$40 [Sea08]. US$65.00 [Cha09]. usability [ZDM10]. usage [ERS04, Rou10b]. use [SH07]. used [PAK01]. User [Ger09, JK09, SK03, ZDM10]. user-assisted [SK03]. user-based [ZDM10]. Using [BG10, Ger04, ASP02, ASR09, BKP05, BLE^{+}08, DST07, Sch07, SLSN10, Ste02, TK04, UKI01, VB02, ZD08].

validation [KKGL10, OEMD08]. validator [NP07]. value [TGH^{+}04]. values [SLSN10]. variability [HBR10]. variable [CDTM10]. version [SCRS08]. Viability [CDN10]. views [GDG06]. Visual [DL09]. visualization [Kos03, KELN10]. Visualizing [FG04]. volatility [FSC^{+}11, ZWP03]. vs [GJ09]. vulnerability [HO05].

WARE [DFT04]. warranty [SZ01]. way [HMG03]. WCRE [HZD10]. web [KM07, BDDT07, BT04, DST07, DFT04, ERS04, LBLA04, Lit04, RDT^{+}07, RPTG08, SH07, TMSM02, TGH^{+}04, TR04, TRPG06, VB02]. Web-based [LBLA04, VB02]. web-localization [KM07]. Web-usage [ERS04]. webmining [ZD08]. well [LFB07]. well-formed [LFB07]. Wiley [Cha09]. Wiley-Interscience [Cha09]. Windows [Ger09]. within [KM02, KM04]. workflows [ACDS02]. Workshops [MY10, RMSL11]. World [Sea08]. WSDL [SH07]. WSDL-Test [SH07]. WSE [BDDT07].

X [Sea08]. XML [LCC^{+}03]. XML-based [LCC^{+}03].

References

Ambrosio:2011:MSS

[ABRF11] Bernardo Giori Ambrósio, José Luis Braga, and Moisés A. Resende-Filho. Modeling and scenario simulation for decision support in management of requirements activities in software projects. Journal of Software Maintenance and Evolution: Research and Practice, 23 (1):35–50, January 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Aversano:2002:AMS

[ACDS02] Lerina Aversano, Gerardo Canfora, Andrea De Lucia, and Silvio Stefanucci. Automating the management of software maintenance workflows in a large software enterprise: a case study. Journal of Software Maintenance and Evolution: Research and Practice, 14 (4):229–255, July 2002. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Andrade:2002:SCC

[AFGK02] L. Andrade, J. L. Fiadeiro, J. Gouveia, and G. Koutsoukos. Separating computation, coordination and configuration. Journal of Software Maintenance and Evolution: Research and Practice, 14(5):353–369, September 2002. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

April:2005:SMM

[AHAD05] Alain April, Jane Huffman Hayes, Alain Abran, and Reiner Dumke. Software Maintenance Maturity Model (SM^{mm}): the software maintenance process model. Journal of Software Maintenance and Evolution: Research and Practice, 17(3): 197–223, May 2005. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Alaghband:2011:RMN

[Ala11] Gita Alaghband. Relationship model: a network model for integrating human expertise with systematic distributed processes. Journal of Software Maintenance and Evolution: Research and Practice, 23 (2):109–135, March 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Adamek:2005:CCE

[AP05] J. Adamek and F. Plasil. Component composition errors and update atomicity: static analysis. Journal of Software Maintenance and Evolution: Research and Practice, 17(5):363– 377, September 2005. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

AlDallal:2005:RCB

[AS05] Jehad Al Dallal and Paul Sorenson. Reusing class-based test cases for testing object-oriented framework interface classes. Journal of Software Maintenance and Evolution: Research and Practice, 17(3):169–196, May 2005. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Ahn:2003:SMP

[ASKK03] Yunsik Ahn, Jungseok Suh, Seungryeol Kim, and Hyunsoo Kim. The software maintenance project effort estimation model based on function points. Journal of Software Maintenance and Evolution: Research and Practice, 15(2):71–85, March 2003. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Abran:2002:FSU

[ASP02] Alain Abran, Ilionar Silva, and Laura Primera. Field studies using functional size measurement in building estimation models for software maintenance. Journal of Software Maintenance and Evolution: Research and Practice, 14 (1):31–64, January 2002. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Anan:2009:ACS

[ASR09] Muhammad Anan, Hossein Saiedian, and Jungwoo Ryoo. An architecture-centric software maintainability assessment using information theory. Journal of Software Maintenance and Evolution: Research and Practice, 21(1):1–18, January 2009. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Aversano:2004:ASI

[AT04] Lerina Aversano and Maria Tortorella. An assessment strategy for identifying legacy system evolution requirements in eBusiness context. Journal of Software Maintenance and Evolution: Research and Practice, 16(4–5):255–276, July/October 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Benestad:2009:USM

[BAA09] Hans Christian Benestad, Bente Anda, and Erik Arisholm. Understanding software maintenance and evolution by analyzing individual changes: a literature review. Journal of Software Maintenance and Evolution: Research and Practice, 21(6):349–378, November 2009. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Balmas:2004:DDG

[Bal04] F. Balmas. Displaying dependence graphs: a hierarchical approach. Journal of Software Maintenance and Evolution: Research and Practice, 16 (3):151–185, May 2004. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Bellini:2005:PDP

[BCG^{+}05] Emilio Bellini, Gerardo Canfora, Félix García, Mario Piattini, and Corrado Aaron Visaggio. Pair designing as practice for enforcing and diffusing design knowledge. Journal of Software Maintenance and Evolution: Research and Practice, 17(6):401–423, November 2005. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Bolchini:2007:ESI

[BDDT07] Davide Bolchini, Thomas Dean, Damiano Distante, and Scott Tilley. Editorials: Special issue on Web Site Evolution (WSE 2006). Journal of Software Maintenance and Evolution: Research and Practice, 19(5):277–279, September 2007. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Bruno:2011:KCE

[BDJ^{+}11] Giorgio Bruno, Frank Dengler, Ben Jennings, Rania Khalaf, Selmin Nurcan, Michael Prilla, Marcello Sarini, Rainer Schmidt, and Rito Silva. Key challenges for enabling agile BPM with social software. Journal of Software Maintenance and Evolution: Research and Practice, 23(4):297–326, June 2011. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Bisbal:2002:CDS

[BG02] Jesús Bisbal and Jane Grimson. Consistent database sampling as a database prototyping approach. Journal of Software Maintenance and Evolution: Research and Practice, 14(6):447–459, November 2002. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Buffett:2010:UCM

[BG10] Scott Buffett and Liqiang Geng.

Using classification methods to label tasks in process mining. Journal of Software Maintenance and Evolution: Research and Practice, 22(6–7):497–517, October/November 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Bergin:2003:CSA

[BK03] Susan Bergin and John Keating. A case study on the adaptive maintenance of an Internet application. Journal of Software Maintenance and Evolution: Research and Practice, 15 (4):245–264, July 2003. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Bril:2005:ASI

[BKP05] R. J. Bril, R. L. Krikhaar, and A. Postma. Architectural support in industry: a reflection using C-POSH. Journal of Software Maintenance and Evolution: Research and Practice, 17 (1):3–25, January 2005. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Barry:2007:HSP

[BKS07] Evelyn J. Barry, Chris F. Kemerer, and Sandra A. Slaughter. How software process automation affects software evolution: a longitudinal empirical analysis. Journal of Software Maintenance and Evolution: Research and Practice, 19 (1):1–31, January 2007. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Black:2001:CRE

[Bla01] Sue Black. Computing ripple effect for software maintenance. Journal of Software Maintenance and Evolution: Research and Practice, 13 (4):263–279, July 2001. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Buckley:2008:ETC

[BLE^{+}08] Jim Buckley, Andrew P. LeGear, Chris Exton, Ross Cadogan, Trevor Johnston, Bill Looby, and Rainer Koschke. Encapsulating targeted component abstractions using software reflexion modelling. Journal of Software Maintenance and Evolution: Research and Practice, 20(2):107–134, March 2008. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Bocco:2005:ACI

[BMP05] Marcela Genero Bocco, Daniel L. Moody, and Mario Piattini. Assessing the capability of internal metrics as early indicators of maintenance effort through experimentation. Journal of Software Maintenance and Evolution: Research and Practice, 17 (3):225–246, May 2005. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Bennett:2008:SET

[BMS^{+}08] C. Bennett, D. Myers, M.-A.

Storey, D. M. German, D. Ouellet, M. Salois, and P. Charland. A survey and evaluation of tool features for understanding reverse-engineered sequence diagrams. Journal of Software Maintenance and Evolution: Research and Practice, 20(4):291–315, July 2008. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Buckley:2005:TTS

[BMZ^{+}05] Jim Buckley, Tom Mens, Matthias Zenger, Awais Rashid, and Günter Kniesel. Towards a taxonomy of software change. Journal of Software Maintenance and Evolution: Research and Practice, 17(5):309–332, September 2005. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Bollin:2008:CLF

[Bol08] Andreas Bollin. Concept location in formal specifications. Journal of Software Maintenance and Evolution: Research and Practice, 20(2):77– 105, March 2008. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Bhatt:2006:IMF

[BSAM06] Pankaj Bhatt, Gautam Shroff, C. Anantaram, and Arun Kumar Misra. An influence model for factors in outsourced software maintenance. Journal of Software Maintenance and Evolution: Research and Practice, 18(6):385–423, November 2006. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Bibi:2010:BBA

[BSGK10] S. Bibi, I. Stamelos, G. Gerolimos, and V. Kollias. BBN based approach for improving the software development process of an SME — a case study. Journal of Software Maintenance and Evolution: Research and Practice, 22(2), March 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Boldyreff:2004:WSE

[BT04] Cornelia Boldyreff and Paolo Tonella. Web site evolution. Journal of Software Maintenance and Evolution: Research and Practice, 16(1–2):1–4, January/April 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Chapin:2001:A

[CC01] Ned Chapin and Aniello Cimitile. Announcement. Journal of Software Maintenance and Evolution: Research and Practice, 13(1):1, January 2001. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Caffery:2010:MSD

[CD10] Fergal Mc Caffery and Alec Dorling. Medi SPICE development. Journal of Software Maintenance and Evolution: Research and Practice, 22 (4):255–268, June 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Capilla:2010:VCD

[CDN10] Rafael Capilla, Juan C. Dueñas, and Francisco Nava. Viability for codifying and documenting architectural design decisions with tool support. Journal of Software Maintenance and Evolution: Research and Practice, 22(2):81–119, March 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Canfora:2011:EPN

[CDR11a] Gerardo Canfora, Darren Dalcher, and David Raffo. Editorials: Preparing for a new era. Journal of Software Maintenance and Evolution: Research and Practice, 23(1):1–2, January 2011. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Canfora:2011:EMM

[CDR^{+}11b] Gerardo Canfora, Darren Dalcher, David Raffo, Victor R. Basili, Juan Fernández-Ramil, Václav Rajlich, Keith Bennett, Liz Burd, Malcolm Munro, Sophia Drossopoulou, Barry Boehm, Susan Eisenbach, Greg Michaelson, Darren Dalcher, Peter Ross, Paul D. Wernick, and Dewayne E. Perry. Editorials: In memory of Manny Lehman, ‘father of software evolution’. Journal of Software Maintenance and Evolution: Research and Practice, 23(3):137–144, April 2011. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Ceccato:2010:MLD

[CDTM10] Mariano Ceccato, Thomas Roy Dean, Paolo Tonella, and Davide Marchignoli. Migrating legacy data structures based on variable overlay to Java. Journal of Software Maintenance and Evolution: Research and Practice, 22(3):211–237, April 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Chapin:2009:BRS

[Cha09] Ned Chapin. Book review: Software Maintenance Management: Evaluation and Continuous Improvement, Authored by Alain April and Alain Abran Wiley-Interscience, Hoboken NJ, U.S.A., and IEEE Computer Society, Los Alamitos CA, U.S.A. Copyright 2008; 314 pages ISBN 978-0470-14707-8 (soft cover) US$65.00. Journal of Software Maintenance and Evolution: Research and Practice, 21(1):73–77, January 2009. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Chapin:2001:TSE

[CHK^{+}01] Ned Chapin, Joanne E. Hale, Khaled Md. Khan, Juan F. Ramil, and Wui-Gee Tan. Types of software evolution and software maintenance. Journal of Software Maintenance and Evolution: Research and Practice, 13(1):3–30, January 2001. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Cook:2006:ESS

[CHLW06] Stephen Cook, Rachel Harrison, Meir M. Lehman, and Paul Wernick. Evolution in software systems: foundations of the SPE classification scheme. Journal of Software Maintenance and Evolution: Research and Practice, 18(1):1–35, January 2006. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Comstock:2011:EDS

[CJD11] Craig Comstock, Zhizhong Jiang, and Jim Davies. Economies and diseconomies of scale in software development. Journal of Software Maintenance and Evolution: Research and Practice, 23(8):533–548, December 2011. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Chaim:2004:DSB

[CMJ04] Marcos L. Chaim, José C. Maldonado, and Mario Jino. A debugging strategy based on the requirements of testing. Journal of Software Maintenance and Evolution: Research and Practice, 16(4–5):277–308, July/October 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Cremer:2002:GBT

[CMW02] Katja Cremer, André Marburger, and Bernhard Westfechtel. Graph-based tools for re-engineering. Journal of Software Maintenance and Evolution: Research and Practice, 14 (4):257–292, July 2002. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Chen:2009:RTS

[CPU09] Yanping Chen, Robert L. Probert, and Hasan Ural. Regression test suite reduction based on SDL models of system requirements. Journal of Software Maintenance and Evolution: Research and Practice, 21(6):379–405, November 2009. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Chung:2003:ABS

[CS03] Lawrence Chung and Nary Subramanian. Architecture-based semantic evolution of embedded remotely controlled systems.

Journal of Software Maintenance and Evolution: Research and Practice, 15(3):145–190, May 2003. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

DiPenta:2008:ESI

[DAH08] Massimiliano Di Penta, Giuliano Antoniol, and Mark Harman. Editorials: Special issue on search-based software maintenance. Journal of Software Maintenance and Evolution: Research and Practice, 20(5):317–319, September 2008. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Daneva:2010:BUC

[Dan10] Maya Daneva. Balancing uncertainty of context in ERP project estimation: an approach and a case study. Journal of Software Maintenance and Evolution: Research and Practice, 22 (5):329–357, August 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

DelRosso:2006:CET

[Del06] Christian Del Rosso. Continuous evolution through software architecture evaluation: a case study. Journal of Software Maintenance and Evolution: Research and Practice, 18(5):351–383, September 2006. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

DelRosso:2009:CAK

[Del09] Christian Del Rosso. Comprehend and analyze knowledge networks to improve software evolution. Journal of Software Maintenance and Evolution: Research and Practice, 21 (3):189–215, May 2009. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

DiLucca:2004:REW

[DFT04] Giuseppe Antonio Di Lucca, Anna Rita Fasolino, and Porfirio Tramontana. Reverse engineering Web applications: the WARE approach. Journal of Software Maintenance and Evolution: Research and Practice, 16(1–2):71–101, January/April 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

DiFrancescomarino:2011:FCS

[DGR^{+}11] Chiara Di Francescomarino, Chiara Ghidini, Marco Rospocher, Luciano Serafini, and Paolo Tonella. A framework for the collaborative specification of semantically annotated business processes. Journal of Software Maintenance and Evolution: Research and Practice, 23 (4):261–295, June 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Desnos:2008:SBM

[DHT^{+}08] Nicolas Desnos, Marianne Huchard, Guy Tremblay, Christelle Urtado, and Sylvain Vauttier. Search-based many-to-one component substitution. Journal of Software Maintenance and Evolution: Research and Practice, 20(5):321–344, September 2008. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Dig:2006:HDA

[DJ06] Danny Dig and Ralph Johnson.

How do APIs evolve? A story of refactoring. Journal of Software Maintenance and Evolution: Research and Practice, 18 (2):83–107, March 2006. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

DAmbros:2009:VSE

[DL09] Marco D’Ambros and Michele Lanza. Visual software evolution reconstruction. Journal of Software Maintenance and Evolution: Research and Practice, 21(3):217–232, May 2009. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

DiPenta:2007:ESI

[DM07] Massimiliano Di Penta and Leon Moonen. Editorials: Special issue on source code analysis and manipulation (SCAM 2006). Journal of Software Maintenance and Evolution: Research and Practice, 19(4): 203–204, July 2007. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Ducasse:2006:ECD

[DNR06] Stéphane Ducasse, Oscar Nierstrasz, and Matthias Rieger. On the effectiveness of clone detection by string matching. Journal of Software Maintenance and Evolution: Research and Practice, 18(1):37–58, January 2006. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Donzelli:2003:TSM

[Don03] Paolo Donzelli. Tailoring the software maintenance process to better support complex systems evolution projects. Journal of Software Maintenance and Evolution: Research and Practice, 15(1):27–40, January 2003. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

DeLucia:2007:ISP

[DST07] Andrea De Lucia, Giuseppe Scanniello, and Genoveffa Tortora. Identifying similar pages in Web applications using a competitive clustering algorithm. Journal of Software Maintenance and Evolution: Research and Practice, 19(5): 281–296, September 2007. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Ducasse:2003:DRE

[DT03] S. Ducasse and S. Tichelaar. Dimensions of reengineering environment infrastructures.

Journal of Software Maintenance and Evolution: Research and Practice, 15(5):345–373, September 2003. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Eierman:2007:PSM

[ED07] Michael A. Eierman and Mark T. Dishaw. The process of software maintenance: a comparison of object-oriented and third-generation development languages. Journal of Software Maintenance and Evolution: Research and Practice, 19(1):33–47, January 2007. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Erwig:2007:PFP

[EFP07] Martin Erwig, Zhe Fu, and Ben Pflaum. Parametric Fortran: program generation in scientific computing. Journal of Software Maintenance and Evolution: Research and Practice, 19 (3):155–182, May 2007. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Erol:2010:CBS

[EGH^{+}10] Selim Erol, Michael Granitzer, Simone Happ, Sami Jantunen, Ben Jennings, Paul Johannesson, Agnes Koschmider, Selmin Nurcan, Davide Rossi, and Rainer Schmidt. Combining BPM and social software: contradiction or chance? Journal of Software Maintenance and Evolution: Research and Practice, 22(6–7):449–476, October/November 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Etzkorn:2002:SEM

[EGJ02] Letha H. Etzkorn, Sampson Gholston, and William E. Hughes Jr. A semantic entropy metric. Journal of Software Maintenance and Evolution: Research and Practice, 14 (4):293–310, July 2002. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

El-Ramly:2004:AWU

[ERS04] Mohammad El-Ramly and Eleni Stroulia. Analysis of Webusage behavior for focused Web sites: a case study. Journal of Software Maintenance and Evolution: Research and Practice, 16(1–2):129–150, January/ April 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Fischer:2004:VFE

[FG04] Michael Fischer and Harald Gall. Visualizing feature evolution of large-scale software based on problem and modification report data. Journal of Software Maintenance and Evolution: Research and Practice, 16(6):385–403, November 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Fogelstrom:2010:IAP

[FGSO10] Nina Dzamashvili Fogelström, Tony Gorschek, Mikael Svahnberg, and Peo Olsson. The impact of agile principles on market-driven software product development. Journal of Software Maintenance and Evolution: Research and Practice, 22 (1):53–80, January 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Ferreira:2011:RRR

[FSC^{+}11] Susan Ferreira, Dan Shunk, James Collofello, Gerald Mackulak, and AmyLou Dueck. Reducing the risk of requirements volatility: findings from an empirical survey. Journal of Software Maintenance and Evolution: Research and Practice, 23 (5):375–393, August 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Garcia-Cabrera:2002:EHS

[GCRFPL02] Lina García-Cabrera, María José Rodríguez-Fórtiz, and José Parets-Llorca. Evolving hypermedia systems: a layered software architecture. Journal of Software Maintenance and Evolution: Research and Practice, 14(5):389–405, September 2002. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Gupta:2010:ECP

[GCS^{+}10] Anita Gupta, Daniela Cruzes, Forrest Shull, Reidar Conradi, Harald Rønneberg, and Einar Landre. An examination of change profiles in reusable and non-reusable software systems. Journal of Software Maintenance and Evolution: Research and Practice, 22 (5):359–380, August 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Gable:2001:LPA

[GCT01] Guy G. Gable, Taizan Chan, and Wui-Gee Tan. Large packaged application software maintenance: a research framework. Journal of Software Maintenance and Evolution: Research and Practice, 13(6): 351–371, November 2001. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Girba:2006:MHA

[GD06] Tudor Gîrba and Stéphane Ducasse. Modeling history to analyze software evolution. Journal of Software Maintenance and Evolution: Research and Practice, 18(3):207–236, May 2006. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Greevy:2006:ASE

[GDG06] Orla Greevy, Stéphane Ducasse, and Tudor Gîrba. Analyzing software evolution through feature views. Journal of Software Maintenance and Evolution: Research and Practice, 18(6):425–456, November 2006. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

German:2004:UST

[Ger04] Daniel M. German. Using software trails to reconstruct the evolution of software. Journal of Software Maintenance and Evolution: Research and Practice, 16(6):367–384, November 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Gerdes:2009:UIM

[Ger09] John Gerdes, Jr. User interface migration of Microsoft Windows applications. Journal of Software Maintenance and Evolution: Research and Practice, 21(3):171–187, May 2009. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Gallagher:2003:GIA

[GHD03] Keith Gallagher, Mark Harman, and Sebastian Danicic. Guaranteed inconsistency avoidance during software evolution. Journal of Software Maintenance and Evolution: Research and Practice, 15(6): 393–416, November 2003. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Gregersen:2009:DUJ

[GJ09] Allan Raundahl Gregersen and Bo Nørregaard Jørgensen. Dynamic update of Java applications — balancing change flexibility vs programming transparency. Journal of Software Maintenance and Evolution: Research and Practice, 21 (2):81–112, March 2009. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Goedicke:2002:PLM

[GZ02] M. Goedicke and U. Zdun. Piecemeal legacy migrating with an architectural pattern language: a case study. Journal of Software Maintenance and Evolution: Research and Practice, 14(1):1–30, January 2002. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Glorie:2009:SLS

[GZvDH09] Marco Glorie, Andy Zaidman, Arie van Deursen, and Lennart Hofland. Splitting a large software repository for easing future software evolution — an industrial experience report. Journal of Software Maintenance and Evolution: Research and Practice, 21(2):113– 141, March 2009. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Hallerbach:2010:CVB

[HBR10] Alena Hallerbach, Thomas Bauer, and Manfred Reichert.

Capturing variability in business process models: the provop approach. Journal of Software Maintenance and Evolution: Research and Practice, 22(6– 7):519–546, October/November 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Heckel:2002:RFR

[HE02] R. Heckel and G. Engels. Relating functional requirements and software architecture: separation and consistency of concerns. Journal of Software Maintenance and Evolution: Research and Practice, 14(5): 371–388, September 2002. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Haapio:2010:SPE

[HE10] Topi Haapio and Anne Eerola.

Software project effort assessment. Journal of Software Maintenance and Evolution: Research and Practice, 22(8): 629–652, December 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Heales:2002:MFA

[Hea02] Jon Heales. A model of factors affecting an information system’s change in state. Journal of Software Maintenance and Evolution: Research and Practice, 14(6):409–427, November 2002. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Higo:2008:MBA

[HKI08] Yoshiki Higo, Shinji Kusumoto, and Katsuro Inoue. A metricbased approach to identifying refactoring opportunities for merging code clones in a Java software system. Journal of Software Maintenance and Evolution: Research and Practice, 20(6):435–461, November 2008. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Heidrich:2010:GOC

[HM10] Jens Heidrich and Jürgen Münch. Goal-oriented customization of software cockpits. Journal of Software Maintenance and Evolution: Research and Practice, 22(5):381– 405, August 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Hayes:2003:OMA

[HMG03] Jane Huffman Hayes, Naresh Mohamed, and Tina Hong Gao. Observe-mine-adopt (OMA): an agile way to enhance software maintainability. Journal of Software Maintenance and Evolution: Research and Practice, 15(5):297–323, September 2003. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Hanebutte:2005:SVM

[HO05] Nadine Hanebutte and Paul W.

Oman. Software vulnerability mitigation as a proper subset of software maintenance. Journal of Software Maintenance and Evolution: Research and Practice, 17(6):379–400, November 2005. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Heston:2011:MQM

[HP11] Keith M. Heston and William Phifer. The multiple quality models paradox: how much ‘best practice’ is just enough? Journal of Software Maintenance and Evolution: Research and Practice, 23(8):517– 531, December 2011. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Hirt:2001:EME

[HS01] Sabine Gabriele Hirt and E. Burton Swanson. Emergent maintenance of ERP: new roles and relationships. Journal of Software Maintenance and Evolution: Research and Practice, 13(6):373–387, November 2001. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Harrison:2002:IHM

[HW02] Matthew S. Harrison and Gwendolyn H. Walton. Identifying high maintenance legacy software. Journal of Software Maintenance and Evolution: Research and Practice, 14(6):429–446, November 2002. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Hassan:2010:EIS

[HZD10] Ahmed E. Hassan, Andy Zaidman, and Massimiliano Di Penta. Editorials: Introduction to the special issue on reverse engineering (WCRE 2008). Journal of Software Maintenance and Evolution: Research and Practice, 22(3): 141–142, April 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Jansen:2006:IDM

[JBBvN06] Slinger Jansen, Gerco Ballintijn, Sjaak Brinkkemper, and Arco van Nieuwland. Integrated development and maintenance for the release, delivery, deployment, and customization of product software: a case study in mass-market ERP software. Journal of Software Maintenance and Evolution: Research and Practice, 18(2): 133–151, March 2006. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Jiang:2008:AAA

[JHHF08] Zhen Ming Jiang, Ahmed E. Hassan, Gilbert Hamann, and Parminder Flora. An automated approach for abstracting execution logs to execution events. Journal of Software Maintenance and Evolution: Research and Practice, 20 (4):249–267, July 2008. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Jiau:2009:AEU

[JK09] Hewijin Christine Jiau and Chia Hung Kao. Assessing the efficacy of user and developer activities in facilitating the development of OSS projects. Journal of Software Maintenance and Evolution: Research and Practice, 21(5):287–314, September 2009. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Jarzabek:2006:UCG

[JL06] Stan Jarzabek and Shubiao Li.

Unifying clones with a generative programming technique: a case study. Journal of Software Maintenance and Evolution: Research and Practice, 18 (4):267–292, July 2006. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Jang:2001:IMS

[JMK01] Y. K. Jang, M. Munro, and Y. R. Kwon. An improved method of selecting regression tests for C++ programs. Journal of Software Maintenance and Evolution: Research and Practice, 13(5):331–350, September 2001. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Jorgensen:2002:IEM

[JS02] Magne Jørgensen and Dag I. K. Sjøberg. Impact of experience on maintenance skills. Journal of Software Maintenance and Evolution: Research and Practice, 14(2):123–146, March 2002. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Jedlitschka:2010:EPM

[JSB10] Andreas Jedlitschka, Outi Salo, and Frank Bomarius. Editorials: Process management. Journal of Software Maintenance and Evolution: Research and Practice, 22(5):325– 328, August 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kanstren:2008:TDU

[Kan08] Teemu Kanstrén. Towards a deeper understanding of test coverage. Journal of Software Maintenance and Evolution: Research and Practice, 20 (1):59–76, January 2008. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kagdi:2007:STA

[KCM07] Huzefa Kagdi, Michael L. Collard, and Jonathan I. Maletic. A survey and taxonomy of approaches for mining software repositories in the context of software evolution. Journal of Software Maintenance and Evolution: Research and Practice, 19(2):77–131, March 2007. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kuhn:2010:SCT

[KELN10] Adrian Kuhn, David Erni, Peter Loretan, and Oscar Nierstrasz. Software cartography: thematic software visualization with consistent layout. Journal of Software Maintenance and Evolution: Research and Practice, 22(3):191–210, April 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kniesel:2005:USE

[KF05] Günter E. Kniesel and Robert E.

Filman. Unanticipated software evolution. Journal of Software Maintenance and Evolution: Research and Practice, 17 (5):307–308, September 2005. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kapser:2006:SAC

[KG06] Cory J. Kapser and Michael W.

Godfrey. Supporting the analysis of clones in software systems. Journal of Software Maintenance and Evolution: Research and Practice, 18(2): 61–82, March 2006. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Khurum:2011:MAE

[KG11] Mahvish Khurum and Tony Gorschek. A method for alignment evaluation of product strategies among stakeholders (MASS) in software intensive product development. Journal of Software Maintenance and Evolution: Research and Practice, 23(7):494–516, November 2011. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kuhne:2010:BPM

[KKGL10] Stefan Kühne, Heiko Kern, Volker Gruhn, and Ralf Laue. Business process modeling with continuous validation. Journal of Software Maintenance and Evolution: Research and Practice, 22(6–7):547–566, October/November 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kakola:2011:ISD

[KKOL11] Timo Käkölä, Mervi Koivulahti-Ojala, and Jani Liimatainen. An information systems design product theory for the class of integrated requirements and release management systems. Journal of Software Maintenance and Evolution: Research and Practice, 23(6): 443–463, October 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kozlov:2008:AMC

[KKSM08] Denis Kozlov, Jussi Koskinen, Markku Sakkinen, and Jouni Markkula. Assessing maintainability change over multiple software releases. Journal of Software Maintenance and Evolution: Research and Practice, 20(1):31–58, January 2008. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kajko-Mattsson:2002:PMM

[KM02] Mira Kajko-Mattsson. Problem management maturity within corrective maintenance. Journal of Software Maintenance and Evolution: Research and Practice, 14(3):197–227, May 2002. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kajko-Mattsson:2004:PWF

[KM04] Mira Kajko-Mattsson. Problems within front-end support. Journal of Software Maintenance and Evolution: Research and Practice, 16(4–5):309–329, July/October 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kagdi:2007:MED

[KM07] Huzefa Kagdi and Jonathan I.

Maletic. Mining evolutionary dependencies from web-localization repositories. Journal of Software Maintenance and Evolution: Research and Practice, 19(5):315–337, September 2007. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kolb:2006:RLC

[KMPY06] Ronny Kolb, Dirk Muthig, Thomas Patzke, and Kazuyuki Yamauchi. Refactoring a legacy component for reuse in a software product line: a case study. Journal of Software Maintenance and Evolution: Research and Practice, 18(2): 109–132, March 2006. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kapur:2008:OSP

[KNTRS08] Puneet Kapur, An Ngo-The, Günther Ruhe, and Andrew Smith. Optimized staffing for product releases and its application at Chartwell Technology. Journal of Software Maintenance and Evolution: Research and Practice, 20(5):365– 386, September 2008. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Koch:2007:SEO

[Koc07] Stefan Koch. Software evolution in open source projects — a large-scale investigation. Journal of Software Maintenance and Evolution: Research and Practice, 19(6):361– 382, November 2007. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kollanus:2011:IMM

[Kol11] Sami Kollanus. ICMM — a maturity model for software inspections. Journal of Software Maintenance and Evolution: Research and Practice, 23 (5):327–341, August 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Koskinen:2002:EEH

[Kos02] Jussi Koskinen. Experimental evaluation of hypertext access structures. Journal of Software Maintenance and Evolution: Research and Practice, 14 (2):83–108, March 2002. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Koschke:2003:SVS

[Kos03] Rainer Koschke. Software visualization in software maintenance, reverse engineering, and re-engineering: a research survey. Journal of Software Maintenance and Evolution: Research and Practice, 15(2):87– 109, March 2003. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kunzle:2011:PTF

[KR11] Vera Künzle and Manfred Reichert. PHILharmonicFlows: towards a framework for objectaware process management. Journal of Software Maintenance and Evolution: Research and Practice, 23(4):205–244, June 2011. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Koskinen:2004:HSI

[KSP04] Jussi Koskinen, Airi Salminen, and Jukka Paakki. Hypertext support for the information needs of software maintainers. Journal of Software Maintenance and Evolution: Research and Practice, 16(3):187– 215, May 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Kontogiannis:2009:ESI

[KTW09] Kostas Kontogiannis, Christos Tjortjis, and Andreas Winter. Editorials: Special issue on the 12th Conference on Software Maintenance and Reengineering (CSMR 2008). Journal of Software Maintenance and Evolution: Research and Practice, 21 (2):79–80, March 2009. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Lavery:2004:MEL

[LBLA04] Janet Lavery, Cornelia Boldyreff, Bin Ling, and Colin Allison. Modelling the evolution of legacy systems to Webbased systems. Journal of Software Maintenance and Evolution: Research and Practice, 16(1–2):5–30, January/ April 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Lu:2003:IDP

[LCC^{+}03] Chih-Wei Lu, William C. Chu, Chih-Hung Chang, Don-Lin Yang, and Wen-Da Lian. Integrating diverse paradigms in evolution and maintenance by an XML-based unified model. Journal of Software Maintenance and Evolution: Research and Practice, 15(3):111–144, May 2003. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Lawrie:2007:ESR

[LFB07] Dawn Lawrie, Henry Feild, and David Binkley. An empirical study of rules for well-formed identifiers. Journal of Software Maintenance and Evolution: Research and Practice, 19 (4):205–229, July 2007. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Li:2009:DRB

[Li09] Wei Li. DynaQoS-RDF: a best effort for QoS-assurance of dynamic reconfiguration of dataflow systems. Journal of Software Maintenance and Evolution: Research and Practice, 21(1):19–48, January 2009. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Light:2001:MIC

[Lig01] Ben Light. The maintenance implications of the customization of ERP software. Journal of Software Maintenance and Evolution: Research and Practice, 13(6):415–429, November 2001. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Litoiu:2004:MWS

[Lit04] Marin Litoiu. Migrating to Web services: a performance engineering approach. Journal of Software Maintenance and Evolution: Research and Practice, 16(1–2):51–70, January/ April 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Lo:2008:MTR

[LKL08] David Lo, Siau-Cheng Khoo, and Chao Liu. Mining temporal rules for software maintenance. Journal of Software Maintenance and Evolution: Research and Practice, 20 (4):227–247, July 2008. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Lee:2011:EMT

[LKP11] Chong Hyung Lee, Yong-Ha Kim, and Dong Ho Park. Evaluation of multi-tasking software system performance with consideration of module dependency. Journal of Software Maintenance and Evolution: Research and Practice, 23 (5):361–374, August 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Lehman:2002:BML

[LKR02] M. M. Lehman, G. Kahen, and J. F. Ramil. Behavioural modelling of long-lived evolution processes — some issues and an example. Journal of Software Maintenance and Evolution: Research and Practice, 14(5):335–351, September 2002. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Lester:2010:IRC

[LWMW10] N. G. Lester, F. G. Wilkie, D. McFall, and M. P. Ware. Investigating the role of CMMI with expanding company size for small- to medium-sized enterprises. Journal of Software Maintenance and Evolution: Research and Practice, 22 (1):17–31, January 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Li:2001:COB

[LYC01] Yang Li, Hongji Yang, and William Chu. A conceptoriented belief revision approach to domain knowledge recovery from source code. Journal of Software Maintenance and Evolution: Research and Practice, 13(1):31–52, January 2001. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Mahdian:2009:RTU

[MAP09] Alireza Mahdian, Anneliese Amschler Andrews, and Orest Jacob Pilskalns. Regression testing with UML software designs: A survey. Journal of Software Maintenance and Evolution: Research and Practice, 21 (4):253–286, July 2009. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

McBride:2010:OTP

[McB10] Tom McBride. Organisational theory perspective on process capability measurement scales. Journal of Software Maintenance and Evolution: Research and Practice, 22(4):243–254, June 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Monteiro:2011:DCI

[MdO11] Luis Felipe Salin Monteiro and Kathia Marçal de Oliveira. Defining a catalog of indicators to support process performance analysis. Journal of Software Maintenance and Evolution: Research and Practice, 23 (6):395–422, October 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Makaroff:2005:DEH

[MH05] Dwight J. Makaroff and Norman C. Hutchinson. Development and evolution of a heterogeneous continuous media server: a case study. Journal of Software Maintenance and Evolution: Research and Practice, 17(2):143–167, March 2005. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Mende:2009:ECS

[MKB09] Thilo Mende, Rainer Koschke, and Felix Beckwermert. An evaluation of code similarity identification for the grow-and-prune model. Journal of Software Maintenance and Evolution: Research and Practice, 21 (2):143–169, March 2009. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Miller:2011:MCT

[MM11a] Keith Miller and Wendy Mac-Caull. Model checking timed properties of healthcare processes. Journal of Software Maintenance and Evolution: Research and Practice, 23(4): 245–260, June 2011. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Mishra:2011:CSP

[MM11b] Deepti Mishra and Alok Mishra.

Complex software project development: agile methods adoption. Journal of Software Maintenance and Evolution: Research and Practice, 23(8):549– 564, December 2011. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Merlo:2003:FFR

[MMD03] Ettore Merlo, Ian McAdam, and Renato De Mori. Feedforward and recurrent neural networks for source code informal information analysis. Journal of Software Maintenance and Evolution: Research and Practice, 15(4):205–244, July 2003. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Memon:2005:ART

[MNX05] Atif Memon, Adithya Nagarajan, and Qing Xie. Automating regression testing for evolving GUI software. Journal of Software Maintenance and Evolution: Research and Practice, 17(1):27–64, January 2005. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Magazinius:2010:SCE

[MPÖ10] Ana Magazinius, Joakim Pernstål, and Peter Öhman. Software cost estimation inhibitors — a case study in automotive context. Journal of Software Maintenance and Evolution: Research and Practice, 22(5):407– 427, August 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Mohamed:2011:MHC

[MRE11] Abdallah Mohamed, Guenther Ruhe, and Armin Eberlein. Mismatch handling for COTS selection: a case study. Journal of Software Maintenance and Evolution: Research and Practice, 23(3):145–178, April 2011. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Mens:2004:AEL

[MRG04] Tom Mens, Juan F. Ramil, and Michael W. Godfrey. Analyzing the evolution of largescale software. Journal of Software Maintenance and Evolution: Research and Practice, 16(6):363–365, November 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Mansour:2011:UBR

[MTN11] Nashat Mansour, Husam Takkoush, and Ali Nehme. UML-based regression testing for OO software. Journal of Software Maintenance and Evolution: Research and Practice, 23(1): 51–68, January 2011. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Mens:2005:FRG

[MVDJ05] Tom Mens, Niels Van Eetvelde, Serge Demeyer, and Dirk Janssens. Formalizing refactorings with graph transformations. Journal of Software Maintenance and Evolution: Research and Practice, 17(4): 247–276, July 2005. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Mens:2002:SCS

[MW02] Tom Mens and Michel Wermelinger. Separation of concerns for software evolution. Journal of Software Maintenance and Evolution: Research and Practice, 14(5):311–315, September 2002. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Ma:2011:SCM

[MWM^{+}11] Yinglong Ma, Haijiang Wu, Xinyu Ma, Beihong Jin, Tao Huang, and Jun Wei. Stable cohesion metrics for evolving ontologies. Journal of Software Maintenance and Evolution: Research and Practice, 23 (5):343–359, August 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Mecella:2010:EIS

[MY10] Massimo Mecella and Jian Yang. Editorials: Introduction to the special issue with best papers from BPM 2008 Workshops. Journal of Software Maintenance and Evolution: Research and Practice, 22(6– 7):431–434, October/November 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Nah:2001:CES

[NFC01] Fiona Fui-Hoon Nah, Silvana Faja, and Teuta Cata. Characteristics of ERP software maintenance: a multiple case study. Journal of Software Maintenance and Evolution: Research and Practice, 13(6): 399–414, November 2001. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Ng:2001:DFE

[Ng01] Celeste See Pui Ng. A decision framework for enterprise resource planning maintenance and upgrade: A client perspective. Journal of Software Maintenance and Evolution: Research and Practice, 13(6):431– 468, November 2001. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Ng:2010:IBC

[NGA10] Janice Ka-Yee Ng, Yann-Gaël Guéhéneuc, and Giuliano Antoniol. Identification of behavioural and creational design motifs through dynamic analysis. Journal of Software Maintenance and Evolution: Research and Practice, 22(8):597– 627, December 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Nikora:2005:AMS

[NM05] Allen P. Nikora and John C.

Munson. An approach to the measurement of software evolution. Journal of Software Maintenance and Evolution: Research and Practice, 17(1):65– 91, January 2005. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Najumudheen:2011:TCA

[NMS11] E. S. F. Najumudheen, Rajib Mall, and Debasis Samanta. Test coverage analysis based on an object-oriented program model. Journal of Software Maintenance and Evolution: Research and Practice, 23(7): 465–493, November 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Noguera:2007:AEA

[NP07] Carlos Noguera and Renaud Pawlak. AVal: an extensible attribute-oriented programming validator for Java. Journal of Software Maintenance and Evolution: Research and Practice, 19(4):253–275, July 2007. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Nevalainen:2010:ESS

[NS10] Risto Nevalainen and Tomas Schweigert. A European scheme for software process improvement manager training and certification. Journal of Software Maintenance and Evolution: Research and Practice, 22 (4):269–277, June 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Ngo:2007:EBR

[NT07] Minh Ngoc Ngo and Hee Beng Kuan Tan. Empiricalbased recovery and maintenance of input error-correction features. Journal of Software Maintenance and Evolution: Research and Practice, 19(6): 419–450, November 2007. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Nielsen:2010:SNS

[NT10] Peter Axel Nielsen and Gitte Tjørnehøj. Social networks in software process improvement. Journal of Software Maintenance and Evolution: Research and Practice, 22(1):33–51, January 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Ohlsson:2001:MFP

[OAW01] Magnus C. Ohlsson, Anneliese Amschler Andrews, and Claes Wohlin. Modelling faultproneness statistically over a sequence of releases: a case study. Journal of Software Maintenance and Evolution: Research and Practice, 13(3):167–199, May 2001. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

OBrien:2004:EBI

[OBS04] Michael P. O’Brien, Jim Buckley, and Teresa M. Shaft. Expectation-based, inferencebased, and bottom-up software comprehension. Journal of Software Maintenance and Evolution: Research and Practice, 16(6):427–447, November 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Olague:2006:ADI

[OELC06] Hector M. Olague, Letha H.

Etzkorn, Wei Li, and Glenn Cox. Assessing design instability in iterative (agile) objectoriented projects. Journal of Software Maintenance and Evolution: Research and Practice, 18(4):237–266, July 2006. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Olague:2008:EVO

[OEMD08] Hector M. Olague, Letha H. Etzkorn, Sherri L. Messimer, and Harry S. Delugach. An empirical validation of objectoriented class complexity metrics and their ability to predict error-prone classes in highly iterative, or agile, software: a case study. Journal of Software Maintenance and Evolution: Research and Practice, 20 (3):171–197, May 2008. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

OKeeffe:2008:SBR

[OÓ08] Mark O’Keeffe and Mel Ó Cinnéide. Search-based refactoring: an empirical study. Journal of Software Maintenance and Evolution: Research and Practice, 20(5):345–364, September 2008. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Orme:2007:IOD

[OYE07] Anthony M. Orme, Haining Yao, and Letha H. Etzkorn. Indicating ontology data quality, stability, and completeness throughout ontology evolution. Journal of Software Maintenance and Evolution: Research and Practice, 19(1):49–75, January 2007. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Popovic:2001:CSM

[PAK01] Miroslav Popović, Branislav Atlagić, and Vladimir Kovačević. Case study: a maintenance practice used with real-time telecommunications software. Journal of Software Maintenance and Evolution: Research and Practice, 13(2):97– 126, March 2001. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Pino:2010:HML

[PBPV10] Francisco J. Pino, Maria Teresa Baldassarre, Mario Piattini, and Giuseppe Visaggio. Harmonizing maturity levels from CMMI-DEV and ISO/IEC 15504. Journal of Software Maintenance and Evolution: Research and Practice, 22(4): 279–296, June 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Parikh:2001:IGP

[PC01] Girish Parikh and Ned Chapin.

Interview with Girish Parikh. Journal of Software Maintenance and Evolution: Research and Practice, 13(3):201–221, May 2001. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Polo:2007:MBA

[PGRP07] Macario Polo, Ignacio García-Rodríguez, and Mario Piattini. An MDA-based approach for database re-engineering. Journal of Software Maintenance and Evolution: Research and Practice, 19(6):383– 417, November 2007. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Power:2004:MSG

[PM04] James F. Power and Brian A. Malloy. A metrics suite for grammar-based software. Journal of Software Maintenance and Evolution: Research and Practice, 16(6):405– 426, November 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Pantsar-Syvaniemi:2006:OED

[PSTN06] Susanna Pantsar-Syväniemi, Jorma Taramaa, and Eila Niemelä. Organizational evolution of digital signal processing software development. Journal of Software Maintenance and Evolution: Research and Practice, 18(4):293–305, July 2006. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Pfleeger:2005:AEL

[PVvV05] Shari Lawrence Pfleeger, Chris Verhoef, and Hans van Vliet. Analyzing the evolution of large-scale software. Journal of Software Maintenance and Evolution: Research and Practice, 17(1):1–2, January 2005. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Quaglini:2010:ICT

[Qua10] Silvana Quaglini. Information and communication technology for process management in healthcare: a contribution to change the culture of blame. Journal of Software Maintenance and Evolution: Research and Practice, 22(6–7):435–448, October/November 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Ravichandar:2008:ICT

[RABT08] Ramya Ravichandar, James D.

Arthur, Shawn A. Bohner, and David P. Tegarden. Improving change tolerance through capabilities-based design: an empirical analysis. Journal of Software Maintenance and Evolution: Research and Practice, 20(2):135–170, March 2008. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Runeson:2003:TPS

[RAH03] Per Runeson, Carina Andersson, and Martin Höst. Test processes in software product evolution — a qualitative survey on the state of practice. Journal of Software Maintenance and Evolution: Research and Practice, 15(1):41–59, January 2003. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Roy:2010:NMF

[RC10] C. K. Roy and J. R. Cordy. Near-miss function clones in open source software: an empirical study. Journal of Software Maintenance and Evolution: Research and Practice, 22 (3):165–189, April 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Robillard:2010:RCC

[RD10] Martin P. Robillard and Barthélémy Dagenais. Recommending change clusters to support software investigation: an empirical study. Journal of Software Maintenance and Evolution: Research and Practice, 22 (3):143–164, April 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Ricca:2007:HDN

[RDT^{+}07] Filippo Ricca, Massimiliano Di Penta, Marco Torchiano, Paolo Tonella, and Mariano Ceccato. How design notations affect the comprehension of Web applications. Journal of Software Maintenance and Evolution: Research and Practice, 19(5):339–359, September 2007. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Rajlich:2006:IIC

[RG06] Václav Rajlich and Tibor Gyimóthy. IEEE International Conference on Software Maintenance (ICSM2005). Journal of Software Maintenance and Evolution: Research and Practice, 18(2):59–60, March 2006. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Raja:2009:MSE

[RHH09] Uzma Raja, David P. Hale, and Joanne E. Hale. Modeling software evolution defects: a time series approach. Journal of Software Maintenance and Evolution: Research and Practice, 21(1):49–71, January 2009. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Rinderle-Ma:2011:ESI

[RMSL11] Stefanie Rinderle-Ma, Shazia Sadiq, and Frank Leymann. Editorials: Special issue on BPM 2009 Workshops. Journal of Software Maintenance and Evolution: Research and Practice, 23(4):203, June 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Rahgozar:2003:ESL

[RO03] Maseud Rahgozar and Farhad Oroumchian. An effective strategy for legacy systems evolution. Journal of Software Maintenance and Evolution: Research and Practice, 15(5):325– 344, September 2003. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Rout:2010:GES

[Rou10a] Terence P. Rout. Guest editorials: Software process improvement and capability determination: selected articles from SPICE 2009. Journal of Software Maintenance and Evolution: Research and Practice, 22 (4):239–241, June 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Rout:2010:SAP

[Rou10b] Terry Rout. Studies on the assessment process: usage of objective evidence in assessing process capability. Journal of Software Maintenance and Evolution: Research and Practice, 22(4):297–305, June 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Ricca:2008:IWS

[RPTG08] Filippo Ricca, Emanuele Pianta, Paolo Tonella, and Christian Girardi. Improving Web site understanding with keyword-based clustering. Journal of Software Maintenance and Evolution: Research and Practice, 20(1):1–29, January 2008. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Rashid:2005:DET

[RS05] Awais Rashid and Peter Sawyer.

A database evolution taxonomy for object-oriented databases. Journal of Software Maintenance and Evolution: Research and Practice, 17(2):93– 141, March 2005. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Rus:2003:PD

[RSL03] Ioana Rus, Carolyn Seaman, and Mikael Lindvall. Process diversity. Journal of Software Maintenance and Evolution: Research and Practice, 15 (1):1–8, January 2003. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Riva:2011:PBA

[RSSX11] Claudio Riva, Petri Selonen, Tarja Systä, and Jianli Xu. A profile-based approach for maintaining software architecture: an industrial experience report. Journal of Software Maintenance and Evolution: Research and Practice, 23 (1):3–20, January 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Raja:2011:CSP

[RT11] Uzma Raja and Marietta J.

Tretter. Classification of software patches: a text mining approach. Journal of Software Maintenance and Evolution: Research and Practice, 23 (2):69–87, March 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Sheldon:2006:MCC

[SC06] Frederick T. Sheldon and Hong Chung. Measuring the complexity of class diagrams in reverse engineering. Journal of Software Maintenance and Evolution: Research and Practice, 18(5):333–350, September 2006. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Schneidewind:2007:ERU

[Sch07] Norman F. Schneidewind. Experience report on using objectoriented design for software maintenance. Journal of Software Maintenance and Evolution: Research and Practice, 19 (3):183–201, May 2007. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Siy:2008:SBA

[SCRS08] Harvey Siy, Parvathi Chundi, Daniel J. Rosenkrantz, and Mahadevan Subramaniam. A segmentation-based approach for temporal analysis of software version repositories. Journal of Software Maintenance and Evolution: Research and Practice, 20(3):199–222, May 2008. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Siebel:2003:LLP

[SCSR03] Nils T. Siebel, Stephen Cook, Manoranjan Satpathy, and Daniel Rodrìguez. Latitudinal and longitudinal process diversity. Journal of Software Maintenance and Evolution: Research and Practice, 15(1):9– 25, January 2003. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Seaman:2001:BRS

[Sea01] Carolyn B. Seaman. Book review: Software Maintenance: Concepts and Practice. Journal of Software Maintenance and Evolution: Research and Practice, 13(2):143–147, March 2001. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Seaman:2003:BRA

[Sea03] Carolyn B. Seaman. Book review: Advances in Software Maintenance Management: Technologies and Solutions. Journal of Software Maintenance and Evolution: Research and Practice, 15(5): 375–377, September 2003. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Seaman:2006:BRM

[Sea06] Carolyn B. Seaman. Book review: Managing Corporate Information Systems Evolution and Maintenance. Journal of Software Maintenance and Evolution: Research and Practice, 18(4):307–309, July 2006. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Seaman:2008:BRS

[Sea08] Carolyn B. Seaman. Book review: Software Maintenance: Concepts and Practice, Authored by Penny Grubb and Armstrong A. Takang World Scientific, New Jersey. Copyright 2003; 349 pages ISBN 981-238-426-X (paperback) US$40. Journal of Software Maintenance and Evolution: Research and Practice, 20(6):463– 466, November 2008. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Simmons:2006:ITF

[SEW^{+}06] Sharon Simmons, Dennis Edwards, Norman Wilde, Josh Homan, and Michael Groble. Industrial tools for the feature location problem: an exploratory study. Journal of Software Maintenance and Evolution: Research and Practice, 18(6):457–474, November 2006. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Sneed:2007:DUW

[SH07] Harry M. Sneed and Shihong Huang. The design and use of WSDL-Test: a tool for testing Web services. Journal of Software Maintenance and Evolution: Research and Practice, 19(5):297–314, September 2007.

CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Sheldon:2002:MMC

[SJC02] Frederick T. Sheldon, Kshamta Jerath, and Hong Chung. Metrics for maintainability of class inheritance hierarchies. Journal of Software Maintenance and Evolution: Research and Practice, 14(3):147–160, May 2002. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Sartipi:2003:UAA

[SK03] Kamran Sartipi and Kostas Kontogiannis. A user-assisted approach to component clustering. Journal of Software Maintenance and Evolution: Research and Practice, 15(4):265– 295, July 2003. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Shatnawi:2010:FSM

[SLSN10] Raed Shatnawi, Wei Li, James Swain, and Tim Newman. Finding software metrics threshold values using ROC curves. Journal of Software Maintenance and Evolution: Research and Practice, 22(1):1–16, January 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Staron:2011:DMS

[SMKN11] Miroslaw Staron, Wilhelm Meding, Göran Karlsson, and Christer Nilsson. Developing measurement systems: an industrial case study. Journal of Software Maintenance and Evolution: Research and Practice, 23(2):89–107, March 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Spinellis:2009:OHF

[Spi09] Diomidis Spinellis. Optimizing header file include directives. Journal of Software Maintenance and Evolution: Research and Practice, 21(4):233–251, July 2009. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Stallinger:2010:III

[SPPV10] Fritz Stallinger, Reinhold Plösch, Gustav Pomberger, and Jan Vollmar. Integrating ISO/IEC 15504 conformant process assessment and organizational reuse enhancement. Journal of Software Maintenance and Evolution: Research and Practice, 22(4):307–324, June 2010. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Stewart:2002:PPD

[Ste02] B. Stewart. Predicting project delivery rates using the Naive-Bayes classifier. Journal of Software Maintenance and Evolution: Research and Practice, 14 (3):161–179, May 2002. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Sahin:2001:PAW

[SZ01] Izzet Sahin and Fatemeh ‘Mariam’ Zahedi. Policy analysis for warranty, maintenance, and upgrade of software systems. Journal of Software Maintenance and Evolution: Research and Practice, 13(6):469– 493, November 2001. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Tonella:2001:IOO

[TA01] Paolo Tonella and Giulio Antoniol. Inference of objectoriented design patterns. Journal of Software Maintenance and Evolution: Research and Practice, 13(5):309–330, September 2001. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Taylor:2003:MEC

[Tay03] M. Taylor. Maintaining electronic commerce systems: the legal issues. Journal of Software Maintenance and Evolution: Research and Practice, 15 (2):61–69, March 2003. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Tilley:2004:BVT

[TGH^{+}04] S. Tilley, J. Gerdes, T. Hamilton, S. Huang, H. Müller, D. Smith, and K. Wong. On the business value and technical challenges of adopting Web services. Journal of Software Maintenance and Evolution: Research and Practice, 16(1–2):31–50, January/ April 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Tan:2001:AEC

[TK01] Hee Beng Kuan Tan and Juan Tiang Kow. An approach for extracting code fragments that implement functionality from source programs. Journal of Software Maintenance and Evolution: Research and Practice, 13(1):53–75, January 2001. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Tahvildar:2004:IDQ

[TK04] Ladan Tahvildar and Kostas Kontogiannis. Improving design quality using meta-pattern transformations: a metricbased approach. Journal of Software Maintenance and Evolution: Research and Practice, 16(4–5):331–361, July/October 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Taylor:2002:MIW

[TMSM02] M. Taylor, J. McWilliam, J. Sheehan, and A. Mulhaney. Maintenance issues in the Web site development process. Journal of Software Maintenance and Evolution: Research and Practice, 14(2):109–122, March 2002. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Tonella:2004:STW

[TR04] Paolo Tonella and Filippo Ricca. Statistical testing of Web applications. Journal of Software Maintenance and Evolution: Research and Practice, 16 (1–2):103–127, January/April 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Tonella:2006:ASA

[TRPG06] Paolo Tonella, Filippo Ricca, Emanuele Pianta, and Christian Girardi. Automatic support for the alignment of multilingual Web sites. Journal of Software Maintenance and Evolution: Research and Practice, 18(3):153–179, May 2006. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Tortorella:2001:ESB

[TV01] Maria Tortorella and Giuseppe Visaggio. Evaluation of a scenario-based reading technique for analysing process components. Journal of Software Maintenance and Evolution: Research and Practice, 13 (3):149–166, May 2001. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Tolfo:2011:AMO

[TWFF11] Cristiano Tolfo, Raul Sidnei Wazlawick, Marcelo Gitirana Gomes Ferreira, and Fernando Antonio Forcellini. Agile methods and organizational culture: reflections about cultural levels. Journal of Software Maintenance and Evolution: Research and Practice, 23 (6):423–441, October 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Tan:2003:AEI

[TZ03] Hee Beng Kuan Tan and Yuan Zhao. Automated elicitation of inclusion dependencies from the source code for database transactions. Journal of Software Maintenance and Evolution: Research and Practice, 15(6):379–392, November 2003. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Uemura:2001:FPA

[UKI01] Takuya Uemura, Shinji Kusumoto, and Katsuro Inoue. Functionpoint analysis using design specifications based on the Unified Modelling Language. Journal of Software Maintenance and Evolution: Research and Practice, 13(4):223–243, July 2001. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Usaola:2001:MTS

[UVG01] Macario Polo Usaola, Mario Piattini Velthuis, and Francisco Ruiz González. MAN-TOOL: a tool for supporting the software maintenance process. Journal of Software Maintenance and Evolution: Research and Practice, 13(2):77– 95, March 2001. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

V:2002:TML

[VB02] Balaji V. and Sangeetha B.

Testing and maintaining delocalized software systems in a multi-site environment using Web-based tools. Journal of Software Maintenance and Evolution: Research and Practice, 14(3):181–195, May 2002. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

vandenBrand:2004:SEC

[vdBCG04] Mark G. J. van den Brand, Gerardo Canfora, and Tibor Gyimothy. Seventh European Conference on Software Maintenance and Reengineering (CSMR 2003). Journal of Software Maintenance and Evolution: Research and Practice, 16(4–5):217–218, July/October 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Veerman:2004:RML

[Vee04] Niels Veerman. Revitalizing modifiability of legacy assets. Journal of Software Maintenance and Evolution: Research and Practice, 16(4–5):219–254, July/October 2004. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

vanGurp:2005:DPS

[vGBB05] Jilles van Gurp, Sjaak Brinkkemper, and Jan Bosch. Design preservation over subsequent releases of a software product: a case study of Baan ERP. Journal of Software Maintenance and Evolution: Research and Practice, 17(4):277–306, July 2005. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Visaggio:2001:ADI

[Vis01] Giuseppe Visaggio. Ageing of a data-intensive legacy system: symptoms and remedies. Journal of Software Maintenance and Evolution: Research and Practice, 13(5):281– 308, September 2001. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Walkinshaw:2008:IDS

[WBHS08] Neil Walkinshaw, Kirill Bogdanov, Mike Holcombe, and Sarah Salahuddin. Improving dynamic software analysis by applying grammar inference principles. Journal of Software Maintenance and Evolution: Research and Practice, 20 (4):269–290, July 2008. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

White:2008:EFR

[WJRR08] Lee White, Khaled Jaber, Brian Robinson, and Václav Rajlich.

Extended firewall for regression testing: an experience report. Journal of Software Maintenance and Evolution: Research and Practice, 20(6):419– 433, November 2008. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Wang:2001:CSR

[WSH01] Shuanglin Wang, Stephen R.

Schach, and Gillian Z. Heller. A case study in repeated maintenance. Journal of Software Maintenance and Evolution: Research and Practice, 13 (2):127–141, March 2001. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Wang:2006:CBA

[WSWM06] Qianxiang Wang, Junrong Shen, Xiaopeng Wang, and Hong Mei. A component-based approach to online software evolution. Journal of Software Maintenance and Evolution: Research and Practice, 18 (3):181–205, May 2006. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Werth:2010:PPB

[WWEL10] Dirk Werth, Philipp Walter, Andreas Emrich, and Peter Loos. Peer-to-peer business processes — conception, construction, tooling. Journal of Software Maintenance and Evolution: Research and Practice, 22(6–7):477–495, October/November 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Ware:2007:APM

[WWS07] M. P. Ware, F. G. Wilkie, and M. Shapcott. The application of product measures in directing software maintenance activity. Journal of Software Maintenance and Evolution: Research and Practice, 19 (2):133–154, March 2007. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Xiong:2011:OSM

[XXN11] Cheng-Jie Xiong, Min Xie, and Szu-Hui Ng. Optimal software maintenance policy considering unavailable time. Journal of Software Maintenance and Evolution: Research and Practice, 23(1):21–33, January 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Yeh:2002:ESI

[YJ02] Dowming Yeh and Jing-Hwa Jeng. An empirical study of the influence of departmentalization and organizational position on software maintenance. Journal of Software Maintenance and Evolution: Research and Practice, 14(1):65–82, January 2002. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Yu:2006:IPM

[Yu06] Liguo Yu. Indirectly predicting the maintenance effort of opensource software. Journal of Software Maintenance and Evolution: Research and Practice, 18(5):311–332, September 2006. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Zaidman:2008:AIK

[ZD08] Andy Zaidman and Serge Demeyer. Automatic identification of key classes in a software system using webmining techniques. Journal of Software Maintenance and Evolution: Research and Practice, 20(6):387–417, November 2008. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Zhao:2010:EIU

[ZDM10] Luyin Zhao, Fadi P. Deek, and James A. McHugh. Exploratory inspection — a userbased learning method for improving open source software usability. Journal of Software Maintenance and Evolution: Research and Practice, 22(8):653–675, December 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Zenger:2005:KES

[Zen05]  Matthias Zenger.  KERIS:

evolving software with extensible modules. Journal of Software Maintenance and Evolution: Research and Practice, 17(5):333–362, September 2005. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Zou:2009:RBP

[ZGFH09] Ying Zou, Jin Guo, King Chun Foo, and Maokeng Hung. Recovering business processes from business applications. Journal of Software Maintenance and Evolution: Research and Practice, 21(5):315–348, September 2009. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Zhang:2011:CBS

[ZHB11] Min Zhang, Tracy Hall, and Nathan Baddoo. Code bad smells: a review of current knowledge. Journal of Software Maintenance and Evolution: Research and Practice, 23 (3):179–202, April 2011. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Zaidman:2008:EIS

[ZHLG08] Andy Zaidman, Abdelwahab Hamou-Lhadj, and Orla Greevy. Editorials: Introduction to the special issue on Program Comprehension through Dynamic Analysis (PCODA). Journal of Software Maintenance and Evolution: Research and Practice, 20(4):223–225, July 2008. CODEN JSMECT.

ISSN 1532-060X (print), 1532-0618 (electronic).

Zhang:2010:BAH

[ZLL^{+}10] Lu Zhang, Jing Luo, He Li, Jiasu Sun, and Hong Mei. A biting-down approach to hierarchical decomposition of object-oriented systems based on structure analysis. Journal of Software Maintenance and Evolution: Research and Practice, 22(8):567–596, December 2010. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Zhang:2007:ACA

[ZR07] Weilei Zhang and Barbara G.

Ryder. Automatic construction of accurate application call graph with library call abstraction for Java. Journal of Software Maintenance and Evolution: Research and Practice, 19 (4):231–252, July 2007. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Zhang:2003:DSV

[ZWP03] Xiaoni Zhang, John Windsor, and Robert Pavur. Determinants of software volatility: a field study. Journal of Software Maintenance and Evolution: Research and Practice, 15 (3):191–204, May 2003. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Zhao:2002:CIA

[ZYXX02] Jianjun Zhao, Hongji Yang, Liming Xiang, and Baowen Xu. Change impact analysis to support architectural evolution. Journal of Software Maintenance and Evolution: Research and Practice, 14(5):317– 333, September 2002. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).